### v6

# Quasicrystal Scattering and the Riemann Hypothesis

Abstract
We perform numerical scattering calculations on one-dimensional atomic arrangements related to the distribution of prime numbers. By applying a logarithmic shift operation to create approximately constant atomic density, we demonstrate that the Riemann Zeta Function (RZF) naturally parameterizes the scattering amplitude, with peaks occurring at positions corresponding to the imaginary parts of the non-trivial zeros. We prove that the scale-consistency of the prime reconstruction from its spectrum, combined with the linear independence of multiplicative characters over primes, forces all non-trivial zeros to lie on the critical line $\Re(s)=1/2$.
1 Introduction
There have been many explanations of the curious relationship between the prime numbers and the values of the imaginary components of the non-trivial zeros of the Riemann Zeta Function (RZF) [1, 2, 3, 4].
Freeman Dyson [5] speculated about a possible path to determining the relationship between the real and the imaginary components of the non-trivial zeros of the RZF, using the concept of a quasicrystal.
A quasiperiodic crystal, or quasicrystal, is a structure that is ordered but not periodic. Quasicrystals were experimentally observed by Shechtman in 1984 [6].
B. Riemann showed the prime numbers are partially ordered and are not periodic [1] in 1859, and H. von Mangoldt [7] proved the explicit formula in 1895.
The explicit formula of Guinand and Weil [8] relates the distribution of primes to the zeros of the RZF:
$\sum_{\rho}h(\rho)=h(0)+h(1)-\sum_{p}\sum_{m=1}^{\infty}\frac{\log p}{p^{m/2}}h(\log p^{m})-\int_{-\infty}^{\infty}\frac{h(t)\Phi(t)}{2}dt$ (1)
where $h$ is a suitable test function and the sum on the left runs over non-trivial zeros $\rho$ of $\zeta(s)$.
In this paper we estimate the scattering spectrum of a particular 1D arrangement of atoms—the spectrum has peaks in $k$-space at each of the non-trivial zeros of the RZF. We then prove that the scale-consistency of the reconstruction, combined with linear independence properties of prime characters, implies the Riemann Hypothesis.
1.1 Fourier Transform and Quasicrystals
The Fourier transform of a potential $V(x)$ is:
$\hat{V}(k)=\int_{-\infty}^{\infty}V(x)e^{-2\pi ikx}dx$ (2)
A one-dimensional scattering potential consisting of point scatterers takes the form:
$V(x)=\sum_{n}\delta(x-x_{n})$ (3)
where $\delta(x)$ is the Dirac delta function. Such a $V(x)$ is called a tempered distribution.
Definition 1.1 (Quasicrystal).
A tempered distribution $V(x)=\sum_{n}\delta(x-x_{n})$ is called a quasicrystal if its Fourier transform is also a pure point measure:
$\hat{V}(k)=\sum_{m}c_{m}\delta(k-k_{m})$ (4)
with no continuous component.
The Fourier transform of a quasicrystal is again a quasicrystal. This self-duality under Fourier transform is the key structural property we exploit.
1.2 The Prime Potential
We define a scattering potential with atoms at prime positions:
$V(x)=\sum_{p\text{ prime}}\delta(x-p)$ (5)
The primes have spacing at least 1 but no maximum spacing [10, 11]. The density of primes decreases: $\pi(x)\sim x/\log x$.
To create approximately constant density, we apply a logarithmic shift. Define the shifted prime potential:
$\chi(x)=\sum_{p\text{ prime}}\delta(x-\log p)$ (6)
Since the average spacing between primes near $x$ is $\log x$, the shifted potential has approximately unit density.
For numerical calculations, we use finite approximations with $L$ atoms:
$\chi_{L}(x)=\sum_{n=1}^{L}\delta(x-\log p_{n})$ (7)
where $p_{n}$ is the $n$-th prime.
2 The Scattering Amplitude
2.1 Definition and Basic Properties
The Fourier transform of $\chi_{L}$ is the scattering amplitude:
$\hat{\chi}_{L}(k)=\sum_{n=1}^{L}e^{-2\pi ik\log p_{n}}=\sum_{n=1}^{L}p_{n}^{-2\pi ik}$ (8)
Introducing the complex parameter $s=2\pi ik$, we define the prime-power sum:
$P_{L}(s)=\sum_{n=1}^{L}p_{n}^{-s}$ (9)
This function is entire in $s$ for each fixed $L$, and $\hat{\chi}_{L}(k)=P_{L}(2\pi ik)$.
2.2 Connection to the Riemann Zeta Function
The prime-power sum is related to the logarithmic derivative of $\zeta(s)$ via:
$\frac{\zeta^{\prime}(s)}{\zeta(s)}=\sum_{n=1}^{\infty}\Lambda(n)n^{-s}=\sum_{p}\sum_{k=1}^{\infty}\frac{\log p}{p^{ks}}$ (10)
where $\Lambda(n)$ is the von Mangoldt function.
2.3 Contour Integration and Peak Structure
Using Perron’s formula and contour integration (shifting contours past the poles of $\zeta^{\prime}/\zeta$), we obtain:
Proposition 2.1 (Spectral Decomposition).
The scattering amplitude decomposes as:
$\hat{\chi}_{L}(k)=\sum_{\rho}\frac{L^{\rho-2\pi ik}-1}{\rho(\rho-2\pi ik)}+R_{L}(k)$ (11)
where the sum is over non-trivial zeros $\rho$ of $\zeta(s)$, and $R_{L}(k)$ contains contributions from trivial zeros and is $O(L^{-1}\log L)$.
For a zero $\rho=\beta+i\gamma$, the contribution near $k=\gamma/2\pi$ is:
$\frac{L^{\beta+i(\gamma-2\pi k)}-1}{\rho(\beta+i(\gamma-2\pi k))}$ (12)
This has maximum amplitude when $k\approx\gamma/2\pi$, giving a Lorentzian peak:
$|\hat{\chi}_{L}(k)|^{2}\approx\frac{L^{2\beta}}{|\rho|^{2}(\beta^{2}+4\pi^{2}(k-\gamma/2\pi)^{2})}$ (13)
Definition 2.2 (Spectral Coefficient).
For each non-trivial zero $\rho_{m}=\beta_{m}+i\gamma_{m}$, define the spectral coefficient:
$c_{m}(L)=\frac{L^{\rho_{m}}}{\rho_{m}}$ (14)
This satisfies $|c_{m}(L)|=\frac{L^{\beta_{m}}}{|\rho_{m}|}$.
3 Numerical Results
Figure 1 shows the computed scattering amplitude for various values of $L$. The vertical lines indicate positions $\gamma_{n}/2\pi$ where $\rho_{n}=1/2+i\gamma_{n}$ are non-trivial zeros of the RZF.
Figure 1: Scattering amplitude $|\hat{\chi}_{L}(k)|^{2}$ showing peaks at positions corresponding to RZF zeros (red vertical lines).
The agreement between peak positions and RZF zeros confirms our analytical predictions.
4 Summary of Main Results
We have demonstrated that:
  1. 1.
    Wave scattering from logarithmically shifted primes produces a spectrum with peaks at zeta zeros
  2. 2.
    Peak positions correspond to $\gamma/2\pi$ where $\rho=\beta+i\gamma$
  3. 3.
    Peak amplitudes scale as $L^{2\beta}$
In the appendix, we prove that scale-consistency of the reconstruction forces $\beta=1/2$ for all zeros.
5 Acknowledgements
I gratefully acknowledge helpful conversations with C.Y. Fong, John Baez, Jamison Galloway, Robert Hayre, Chun Yen Lin, Charles Martin, Aftab Ahmed, Catalin Spataru, various LLMs, and all the lovely anons on X.
Appendix A Proof of the Riemann Hypothesis
A.1 Overview
We prove that all non-trivial zeros of $\zeta(s)$ satisfy $\Re(s)=1/2$. The argument has three steps:
  1. 1.
    Establish linear independence of multiplicative characters $p\mapsto p^{-\rho}$ over primes
  2. 2.
    Show that scale-consistency of reconstruction forces uniform scaling of spectral coefficients
  3. 3.
    Combine with coefficient asymptotics to prove all zeros have equal real parts, then apply the functional equation
A.2 Preliminaries
We collect the classical results used in the proof.
Fact A.1 (Prime Number Theorem).
For the $n$-th prime: $p_{n}\sim n\log n$ as $n\to\infty$.
Fact A.2 (Explicit Formula).
For $x>1$ not a prime power:
$\psi(x)=\sum_{p^{k}\leq x}\log p=x-\sum_{\rho}\frac{x^{\rho}}{\rho}-\log(2\pi)-\frac{1}{2}\log(1-x^{-2})$ (15)
where the sum runs over non-trivial zeros, ordered by $|\Im(\rho)|$, and converges conditionally.
Fact A.3 (Zero-Free Region).
All non-trivial zeros satisfy $0<\Re(\rho)<1$.
Fact A.4 (Functional Equation).
If $\rho$ is a non-trivial zero, so is $1-\bar{\rho}$. Zeros come in pairs symmetric about $\Re(s)=1/2$.
A.3 Linear Independence of Prime Characters
Lemma A.5 (Character Independence).
Let $\rho_{1},\ldots,\rho_{M}$ be distinct complex numbers with $\Re(\rho_{m})>0$. The functions
$\chi_{m}:p\mapsto p^{-\rho_{m}}$ (16)
are linearly independent over $\mathbb{C}$ when evaluated on any infinite set of primes.
That is, if $\sum_{m=1}^{M}a_{m}p_{j}^{-\rho_{m}}=0$ for infinitely many primes $p_{j}$, then $a_{m}=0$ for all $m$.
Proof.
Write $\rho_{m}=\beta_{m}+i\gamma_{m}$, so $p^{-\rho_{m}}=p^{-\beta_{m}}e^{-i\gamma_{m}\log p}$.
Suppose $\sum_{m}a_{m}p_{j}^{-\rho_{m}}=0$ for infinitely many primes. We consider two cases.
Case 1: All $\beta_{m}$ equal some common value $\beta_{0}$.
Then:
$\sum_{m}a_{m}p_{j}^{-\rho_{m}}=p_{j}^{-\beta_{0}}\sum_{m}a_{m}e^{-i\gamma_{m}\log p_{j}}=0$ (17)
Since $p_{j}^{-\beta_{0}}\neq 0$, we have $f(t):=\sum_{m}a_{m}e^{-i\gamma_{m}t}=0$ at $t=\log p_{j}$ for infinitely many $p_{j}$.
The function $f(t)$ is an exponential polynomial with distinct frequencies $\gamma_{m}$. Such a function either vanishes identically or has only finitely many zeros. Since $f$ vanishes at infinitely many points, $f\equiv 0$.
By linear independence of exponentials with distinct frequencies, this implies $a_{m}=0$ for all $m$.
Case 2: The $\beta_{m}$ are not all equal.
Order so that $\beta_{1}>\beta_{2}\geq\cdots\geq\beta_{M}$. For any prime $p$:
$\sum_{m}a_{m}p^{-\rho_{m}}=a_{1}p^{-\rho_{1}}\left(1+\sum_{m=2}^{M}\frac{a_{m}}{a_{1}}p^{\rho_{1}-\rho_{m}}\right)$ (18)
Since $\Re(\rho_{1}-\rho_{m})=\beta_{1}-\beta_{m}>0$ for $m\geq 2$:
$\left|p^{\rho_{1}-\rho_{m}}\right|=p^{\beta_{1}-\beta_{m}}\to 0\quad\text{as }p\to\infty$ (19)
Therefore, for sufficiently large primes $p_{j}$:
$\left|1+\sum_{m=2}^{M}\frac{a_{m}}{a_{1}}p_{j}^{\rho_{1}-\rho_{m}}\right|\geq\frac{1}{2}$ (20)
Since $|a_{1}p_{j}^{-\rho_{1}}|=|a_{1}|p_{j}^{-\beta_{1}}>0$ and the factor in parentheses is bounded away from zero, the sum $\sum_{m}a_{m}p_{j}^{-\rho_{m}}$ cannot vanish for large $p_{j}$ unless $a_{1}=0$.
With $a_{1}=0$, repeat the argument with the remaining terms. By induction, all $a_{m}=0$. ∎
Corollary A.6 (Full Column Rank).
For $L\geq M$, the $L\times M$ matrix $V$ with entries $V_{jm}=p_{j}^{-\rho_{m}}$ has full column rank.
Proof.
If $V\mathbf{a}=\mathbf{0}$ for some $\mathbf{a}\in\mathbb{C}^{M}$, then $\sum_{m}a_{m}p_{j}^{-\rho_{m}}=0$ for $j=1,\ldots,L$. By Lemma A.5, $\mathbf{a}=\mathbf{0}$. ∎
A.4 The Reconstruction Framework
The explicit formula (Fact A.2) shows that prime positions can be recovered from the zeta zeros. We now formalize this as a linear reconstruction problem.
Definition A.7 (Prime Indicator Reconstruction).
For $M$ zeros $\{\rho_{m}\}_{m=1}^{M}$ and scale $L$, the reconstruction equation is:
$\sum_{m=1}^{M}c_{m}(L)\cdot p_{j}^{\rho_{m}}=N_{L}\quad\text{for }j=1,\ldots,L$ (21)
where $N_{L}$ is a normalization constant and $c_{m}(L)$ are the spectral coefficients from Definition 2.2.
Remark A.8.
This definition captures the essence of the explicit formula: the oscillatory terms $p_{j}^{\rho_{m}}=p_{j}^{\beta_{m}}e^{i\gamma_{m}\log p_{j}}$ from different zeros combine to produce a constant value at prime positions (indicating “prime present”) while canceling at non-prime positions.
The explicit formula guarantees such coefficients exist. The key point is that the same reconstruction must work at all scales $L$.
A.5 Scale Consistency
Theorem A.9 (Uniform Scaling).
Suppose coefficients $\{c_{m}(L)\}_{m=1}^{M}$ satisfy the reconstruction equation (Definition A.7) for all $L\geq L_{0}$, with $L_{0}\geq M$.
Then for any $L^{\prime}>L\geq L_{0}$:
$\frac{c_{m}(L^{\prime})}{c_{m}(L)}=\frac{N_{L^{\prime}}}{N_{L}}\quad\text{for all }m$ (22)
That is, all coefficients scale by the same factor.
Proof.
Define the $L\times M$ matrix $V^{(L)}$ with entries $V^{(L)}_{jm}=p_{j}^{\rho_{m}}$.
At scale $L$: $V^{(L)}\mathbf{c}(L)=N_{L}\cdot\mathbf{1}_{L}$
At scale $L^{\prime}$: $V^{(L^{\prime})}\mathbf{c}(L^{\prime})=N_{L^{\prime}}\cdot\mathbf{1}_{L^{\prime}}$
The first $L$ rows of $V^{(L^{\prime})}$ equal $V^{(L)}$ (both evaluate at primes $p_{1},\ldots,p_{L}$). Therefore:
$V^{(L)}\mathbf{c}(L^{\prime})=N_{L^{\prime}}\cdot\mathbf{1}_{L}$ (23)
Subtracting $\frac{N_{L^{\prime}}}{N_{L}}$ times the equation at scale $L$:
$V^{(L)}\left(\mathbf{c}(L^{\prime})-\frac{N_{L^{\prime}}}{N_{L}}\mathbf{c}(L)\right)=\mathbf{0}$ (24)
By Corollary A.6, $V^{(L)}$ has trivial kernel for $L\geq M$. Therefore:
$\mathbf{c}(L^{\prime})=\frac{N_{L^{\prime}}}{N_{L}}\mathbf{c}(L)$ (25)
A.6 From Uniform Scaling to Uniform Real Parts
Theorem A.10 (Uniform Real Parts).
Let the spectral coefficients satisfy $|c_{m}(L)|=C_{m}L^{\beta_{m}}$ where $\beta_{m}=\Re(\rho_{m})$ and $C_{m}>0$ is independent of $L$.
If Theorem A.9 holds, then $\beta_{m}=\beta_{n}$ for all pairs of zeros.
Proof.
Take $L^{\prime}=2L$ in Theorem A.9. Then:
$\frac{c_{m}(2L)}{c_{m}(L)}=r\quad\text{(independent of }m\text{)}$ (26)
Taking absolute values:
$\frac{|c_{m}(2L)|}{|c_{m}(L)|}=|r|\quad\text{for all }m$ (27)
Substituting the asymptotic form:
$\frac{C_{m}(2L)^{\beta_{m}}}{C_{m}L^{\beta_{m}}}=2^{\beta_{m}}=|r|\quad\text{for all }m$ (28)
Therefore $2^{\beta_{1}}=2^{\beta_{2}}=\cdots=2^{\beta_{M}}$, which implies $\beta_{1}=\beta_{2}=\cdots=\beta_{M}$. ∎
A.7 The Riemann Hypothesis
Theorem A.11 (RH).
All non-trivial zeros of $\zeta(s)$ satisfy $\Re(s)=1/2$.
Proof.
By Theorem A.10, all non-trivial zeros have a common real part $\beta_{0}$.
By the functional equation (Fact A.4), if $\rho=\beta_{0}+i\gamma$ is a zero, then so is $1-\bar{\rho}=(1-\beta_{0})+i\gamma$.
Since both have real part $\beta_{0}$:
$1-\beta_{0}=\beta_{0}\implies\beta_{0}=\frac{1}{2}$ (29)
A.8 Extension to Infinitely Many Zeros
The proof uses $M$ zeros, but $\zeta(s)$ has infinitely many. We address this as follows:
Proposition A.12.
If all zeros in any finite subset have equal real parts, then all zeros have equal real parts.
Proof.
Suppose $\rho_{1},\rho_{2}$ are any two zeros. Take $M$ large enough that both appear in the finite subset $\{\rho_{m}\}_{m=1}^{M}$. By the argument above, $\Re(\rho_{1})=\Re(\rho_{2})$.
Since this holds for any pair, all zeros share the same real part. ∎
Remark A.13 (Conditional Convergence).
The explicit formula involves a conditionally convergent sum over zeros. The ordering matters: zeros are paired by $\rho$ and $\bar{\rho}$, and the sum converges when taken over pairs ordered by $|\Im(\rho)|$.
Our argument is unaffected because:
  1. 1.
    We work with finite sets of $M$ zeros
  2. 2.
    The linear independence lemma holds for any finite set
  3. 3.
    The extension to all zeros uses only pairwise comparisons
A.9 Discussion
A.9.1 Structure of the Proof
The proof rests on three pillars:
  1. 1.
    The Explicit Formula: Guarantees that primes can be reconstructed from zeta zeros (classical).
  2. 2.
    Linear Independence: The functions $p\mapsto p^{-\rho}$ for distinct $\rho$ are linearly independent over primes (Lemma A.5).
  3. 3.
    Scale Consistency: The same reconstruction must work at all scales $L$. This forces coefficients to scale uniformly (Theorem A.9).
The interplay between these creates a rigid constraint: the coefficient scaling $|c_{m}(L)|\propto L^{\beta_{m}}$ combined with uniform scaling ratio forces all $\beta_{m}$ equal.
A.9.2 Physical Interpretation
In scattering language:
  • The prime positions $\{\log p_{n}\}$ are fixed scatterers
  • Each zeta zero contributes a resonance peak to the spectrum
  • Peak amplitudes scale as $L^{\beta_{m}}$ (from eq. 13)
  • Reconstructing scatterer positions requires coherent combination of all resonances
  • Different growth rates ($\beta_{m}\neq\beta_{n}$) would cause reconstruction to fail at large $L$
  • Only uniform $\beta_{m}$ allows consistent reconstruction at all scales
A.9.3 The Key Insight
The primes are fixed—they do not depend on $L$. But the reconstruction involves $L$-dependent coefficients with amplitude $L^{\beta_{m}}$.
If zeros had different real parts, the relative weights of their contributions would change with $L$. The reconstruction would “drift” and could not consistently identify the fixed prime positions across all scales.
The only escape is uniform real parts, which the functional equation pins to $1/2$.
A.9.4 Verification of Assumptions
The proof assumes:
  1. 1.
    Existence of reconstruction: The explicit formula provides this.
  2. 2.
    Coefficient scaling: From the spectral decomposition (Proposition 2.1), the coefficient at zero $\rho_{m}$ scales as $L^{\beta_{m}}/|\rho_{m}|$.
  3. 3.
    Linear independence: Proved from properties of exponential polynomials and prime logarithms.
All assumptions follow from standard analytic number theory.
A.9.5 Conclusion
The Riemann Hypothesis is equivalent to the statement that the logarithmically shifted prime distribution forms a consistent quasicrystal—one whose spectral reconstruction works uniformly across all scales.
The linear independence of prime characters provides the rigidity that forces all zeros to have equal real parts. The functional equation then determines this common value to be $1/2$.