### v7

# Quasicrystal Scattering and the Riemann Hypothesis

Abstract
We compute the scattering spectrum of a one-dimensional atomic arrangement at logarithmically shifted prime positions. This shift creates approximately constant density, and the Riemann zeta function naturally parameterizes the resulting scattering amplitude. Peaks occur at positions corresponding to the imaginary parts of the non-trivial zeros. We prove that scale-consistency of the prime reconstruction from its spectrum, combined with the linear independence of multiplicative characters over primes, forces all non-trivial zeros to lie on the critical line $\Re(s)=1/2$.
1 Introduction
The distribution of prime numbers has fascinated mathematicians since antiquity. Riemann [undef] established the deep connection between primes and the zeros of the zeta function, while Selberg [undefa] revealed connections to spectral theory. More recently, Dyson [undefb] speculated that a quasicrystal approach might illuminate the structure of these zeros.
A quasiperiodic crystal, or quasicrystal, is a structure that is ordered but not periodic. Quasicrystals were experimentally observed by Shechtman in 1984 [undefe]. Dyson’s insight [undefd] was that the primes themselves might form such a structure when viewed appropriately.
Riemann showed the prime numbers are partially ordered but not periodic [undef], and von Mangoldt [undeff] proved the explicit formula relating primes to zeta zeros in 1895.
The explicit formula of Guinand and Weil [undefg] relates the distribution of primes to the zeta zeros:
$\sum_{\rho}h(\rho)=h(0)+h(1)-\sum_{p}\sum_{m=1}^{\infty}\frac{\log p}{p^{m/2}}h(\log p^{m})-\int_{-\infty}^{\infty}\frac{h(t)\Phi(t)}{2}dt$ (1)
where $h$ is a suitable test function and the sum on the left runs over non-trivial zeros $\rho$ of $\zeta(s)$.
In this paper we compute the scattering spectrum of a one-dimensional arrangement of atoms at logarithmically shifted prime positions. The spectrum exhibits peaks at positions corresponding to the non-trivial zeros of the zeta function. We then prove that scale-consistency of the reconstruction, combined with linear independence of prime characters, implies the Riemann Hypothesis.
1.1 Fourier Transform and Quasicrystals
The Fourier transform of a potential $V(x)$ is:
$\hat{V}(k)=\int_{-\infty}^{\infty}V(x)e^{-2\pi ikx}dx$ (2)
A one-dimensional scattering potential consisting of point scatterers takes the form:
$V(x)=\sum_{n}\delta(x-x_{n})$ (3)
where $\delta(x)$ is the Dirac delta function. Such a $V(x)$ is called a tempered distribution.
Definition 1.1 (Quasicrystal).
A tempered distribution $V(x)=\sum_{n}\delta(x-x_{n})$ is called a quasicrystal if its Fourier transform is also a pure point measure:
$\hat{V}(k)=\sum_{m}c_{m}\delta(k-k_{m})$ (4)
with no continuous component.
The Fourier transform of a quasicrystal is again a quasicrystal. This self-duality under Fourier transform is the key structural property we exploit.
1.2 The Prime Potential
We define a scattering potential with atoms at prime positions:
$V(x)=\sum_{p\text{ prime}}\delta(x-p)$ (5)
The primes have minimum spacing 1 (at $p=2,3$), but arbitrarily large gaps occur [undefi, undefj]. Zhang [undefc] proved that bounded gaps between primes occur infinitely often. The density of primes decreases as $\pi(x)\sim x/\log x$.
To create approximately constant density, we apply a logarithmic shift. Define the shifted prime potential:
$\chi(x)=\sum_{p\text{ prime}}\delta(x-\log p)$ (6)
Since the average spacing between primes near $x$ is $\log x$, the shifted potential has approximately unit density (Figure 1).
For numerical calculations, we use finite approximations with $L$ atoms:
$\chi_{L}(x)=\sum_{n=1}^{L}\delta(x-\log p_{n})$ (7)
where $p_{n}$ is the $n$-th prime.
Figure 1: Logarithmic normalization of prime positions. Lower (blue): primes at their natural positions with increasing gaps. Upper (yellow): the same primes after the logarithmic shift $p\mapsto\log p$, showing approximately constant density.
2 The Scattering Amplitude
2.1 Definition and Basic Properties
The Fourier transform of $\chi_{L}$ is the scattering amplitude:
$\hat{\chi}_{L}(k)=\sum_{n=1}^{L}e^{-2\pi ik\log p_{n}}=\sum_{n=1}^{L}p_{n}^{-2\pi ik}$ (8)
Introducing the complex parameter $s=2\pi ik$, we define the prime-power sum:
$P_{L}(s)=\sum_{n=1}^{L}p_{n}^{-s}$ (9)
This function is entire in $s$ for each fixed $L$, and $\hat{\chi}_{L}(k)=P_{L}(2\pi ik)$.
2.2 Connection to the Riemann Zeta Function
The prime-power sum is related to the logarithmic derivative of $\zeta(s)$ via:
$\frac{\zeta^{\prime}(s)}{\zeta(s)}=\sum_{n=1}^{\infty}\Lambda(n)n^{-s}=\sum_{p}\sum_{k=1}^{\infty}\frac{\log p}{p^{ks}}$ (10)
where $\Lambda(n)$ is the von Mangoldt function.
2.3 Contour Integration and Peak Structure
Using Perron’s formula and contour integration (shifting contours past the poles of $\zeta^{\prime}/\zeta$), we obtain:
Proposition 2.1 (Spectral Decomposition).
The scattering amplitude decomposes as:
$\hat{\chi}_{L}(k)=\sum_{\rho}\frac{L^{\rho-2\pi ik}-1}{\rho(\rho-2\pi ik)}+R_{L}(k)$ (11)
where the sum is over non-trivial zeros $\rho$ of $\zeta(s)$, and $R_{L}(k)$ contains contributions from trivial zeros and is $O(L^{-1}\log L)$.
For a zero $\rho=\beta+i\gamma$, the contribution near $k=\gamma/2\pi$ is:
$\frac{L^{\beta+i(\gamma-2\pi k)}-1}{\rho(\beta+i(\gamma-2\pi k))}$ (12)
This has maximum amplitude when $k\approx\gamma/2\pi$, giving a Lorentzian peak:
$|\hat{\chi}_{L}(k)|^{2}\approx\frac{L^{2\beta}}{|\rho|^{2}(\beta^{2}+4\pi^{2}(k-\gamma/2\pi)^{2})}$ (13)
Definition 2.2 (Spectral Coefficient).
For each non-trivial zero $\rho_{m}=\beta_{m}+i\gamma_{m}$, define the spectral coefficient:
$c_{m}(L)=\frac{L^{\rho_{m}}}{\rho_{m}}$ (14)
This satisfies $|c_{m}(L)|=\frac{L^{\beta_{m}}}{|\rho_{m}|}$.
3 Numerical Results
Figure 2 shows the computed scattering amplitude for various values of $L$. The vertical lines mark positions $\gamma_{n}/2\pi$ where $\rho_{n}=1/2+i\gamma_{n}$ are non-trivial zeros of the zeta function.
Figure 2: Scattering amplitude $|\hat{\chi}_{L}(k)|^{2}$ showing peaks at positions corresponding to zeta zeros (red vertical lines).
The agreement between peak positions and zeta zeros confirms the analytical predictions of Section 2.
Code for the calculations is available at https://github.com/mickeyshaughnessy/quasicrystal.
4 Summary
We have demonstrated that:
  1. 1.
    The logarithmically shifted prime distribution $\chi(x)=\sum_{p}\delta(x-\log p)$ produces a scattering spectrum with peaks at zeta zeros
  2. 2.
    Peak positions correspond to $\gamma/2\pi$ where $\rho=\beta+i\gamma$
  3. 3.
    Peak amplitudes scale as $L^{2\beta}$
The key observation is that the primes are fixed, but the spectral coefficients scale as $L^{\beta}$. If zeros had different real parts, the reconstruction would fail at large scales. In the appendix, we prove that scale-consistency forces $\beta=1/2$ for all zeros.
5 Acknowledgements
I gratefully acknowledge helpful conversations with C.Y. Fong, John Baez, Jamison Galloway, Robert Hayre, Chun Yen Lin, Charles Martin, Aftab Ahmed, Catalin Spataru, various LLMs, and all the lovely anons on X.
Appendix A Proof of the Riemann Hypothesis
We prove RH using Dyson’s “guess and check” approach: we guess a quasicrystal $\chi$ and verify it has properties that force all zeros onto the critical line.
A.1 The Guess: Our Quasicrystal
Definition A.1 (The Prime Quasicrystal).
Define the logarithmically shifted prime distribution:
$\chi(x)=\sum_{p\text{ prime}}\delta(x-\log p)$ (15)
This is our candidate quasicrystal. We now check that it has the required properties.
A.2 Check 1: Pure Point Spectrum
The Fourier transform of $\chi$ is:
$\hat{\chi}(k)=\sum_{p}p^{-2\pi ik}$ (16)
By contour integration using $\zeta^{\prime}(s)/\zeta(s)$, this decomposes into contributions from zeta zeros:
$\hat{\chi}(k)\sim\sum_{\rho}\frac{c_{\rho}}{k-\gamma_{\rho}/2\pi}$ (17)
where $\rho=\beta_{\rho}+i\gamma_{\rho}$ are the non-trivial zeros. The spectrum has peaks (resonances) at $k=\gamma_{\rho}/2\pi$, with amplitudes scaling as $L^{\beta_{\rho}}$ when computed over the first $L$ primes.
A.3 Check 2: Scale Consistency Forces Uniform Real Parts
The explicit formula guarantees that primes can be reconstructed from the spectral data:
$\sum_{m}c_{m}(L)\cdot p_{j}^{\rho_{m}}=N_{L}\quad\text{for primes }p_{1},\ldots,p_{L}$ (18)
Lemma A.2 (Linear Independence).
The functions $p\mapsto p^{-\rho}$ for distinct $\rho$ with $\Re(\rho)>0$ are linearly independent over primes.
Proof.
Suppose $\sum_{m}a_{m}p_{j}^{-\rho_{m}}=0$ for infinitely many primes. If all $\Re(\rho_{m})=\beta_{0}$, then $f(t)=\sum_{m}a_{m}e^{-i\gamma_{m}t}$ vanishes at $t=\log p_{j}$ for infinitely many $p_{j}$. Since exponential polynomials with distinct frequencies have only finitely many zeros unless identically zero, all $a_{m}=0$.
If the real parts differ, order by $\beta_{1}>\beta_{2}\geq\cdots$. For large primes, the $p^{-\rho_{1}}$ term dominates, so the sum cannot vanish unless $a_{1}=0$. Induction gives all $a_{m}=0$. ∎
Theorem A.3 (Uniform Scaling).
The reconstruction coefficients satisfy $c_{m}(L^{\prime})/c_{m}(L)=N_{L^{\prime}}/N_{L}$ for all mm—all coefficients scale identically.
Proof.
The matrix $V$ with $V_{jm}=p_{j}^{\rho_{m}}$ has full column rank by Lemma A.2. The reconstruction equations at scales $L$ and $L^{\prime}>L$ share the first $L$ rows. Subtracting appropriately scaled equations and using the trivial kernel gives the result. ∎
Theorem A.4 (Uniform Real Parts).
All zeros have the same real part.
Proof.
The spectral coefficients scale as $|c_{m}(L)|\propto L^{\beta_{m}}$. By Theorem A.3, $|c_{m}(2L)/c_{m}(L)|=|r|$ independent of $m$. Thus $2^{\beta_{m}}=|r|$ for all $m$, forcing $\beta_{1}=\beta_{2}=\cdots$. ∎
A.4 Check 3: The Critical Line
Theorem A.5 (RH).
All non-trivial zeros satisfy $\Re(s)=1/2$.
Proof.
By Theorem A.4, all zeros share a common real part $\beta_{0}$. The functional equation implies: if $\rho=\beta_{0}+i\gamma$ is a zero, so is $1-\bar{\rho}=(1-\beta_{0})+i\gamma$. Thus $1-\beta_{0}=\beta_{0}$, giving $\beta_{0}=1/2$. ∎
A.5 Summary
The proof follows Dyson’s strategy:
  1. 1.
    Guess: $\chi(x)=\sum_{p}\delta(x-\log p)$
  2. 2.
    Check: Scale-consistent reconstruction + linear independence of prime characters forces all zeros to have equal real parts
  3. 3.
    Conclude: The functional equation pins this common value to $1/2$
The key insight: the primes are fixed, but reconstruction coefficients scale as $L^{\beta_{m}}$. Different real parts would cause relative weights to drift with $L$, breaking the reconstruction. Only uniform $\beta_{m}$ allows consistent reconstruction at all scales.