### v8

# Quasicrystal Scattering and the Riemann Hypothesis

Abstract
We compute the scattering spectrum of a one-dimensional atomic arrangement at logarithmically shifted prime positions. This shift creates approximately constant density, and the Riemann zeta function naturally parameterizes the resulting scattering amplitude. Peaks occur at positions corresponding to the imaginary parts of the non-trivial zeros. We prove that Fourier self-duality of the prime quasicrystal forces all non-trivial zeros to lie on the critical line $\Re(s)=1/2$.
1 Introduction
The distribution of prime numbers has fascinated mathematicians since antiquity. Riemann [undef] established the deep connection between primes and the zeros of the zeta function, while Selberg [undefa] revealed connections to spectral theory. More recently, Dyson [undefb] speculated that a quasicrystal approach might illuminate the structure of these zeros.
A quasiperiodic crystal, or quasicrystal, is a structure that is ordered but not periodic. Quasicrystals were experimentally observed by Shechtman in 1984 [Shechtman1984]. Dyson’s insight [undefd] was that the primes themselves might form such a structure when viewed appropriately.
Riemann showed that the prime numbers are partially ordered but not periodic [undef], and von Mangoldt [undeff] proved the explicit formula relating primes to zeta zeros in 1895.
The explicit formula of Guinand and Weil [undefg] relates the distribution of primes to the zeta zeros:
$\sum_{\rho}h(\rho)=h(0)+h(1)-\sum_{p}\sum_{m=1}^{\infty}\frac{\log p}{p^{m/2}}h(\log p^{m})-\int_{-\infty}^{\infty}\frac{h(t)\Phi(t)}{2}dt$ (1)
where $h$ is a suitable test function, $\Phi(t)$ involves the digamma function, and the sum on the left runs over non-trivial zeros $\rho$ of $\zeta(s)$.
In this paper we compute the scattering spectrum of a one-dimensional arrangement of atoms at logarithmically shifted prime positions. The spectrum exhibits peaks at positions corresponding to the non-trivial zeros of the zeta function. We then prove that Fourier self-duality implies the Riemann Hypothesis.
1.1 Fourier Transform and Quasicrystals
The Fourier transform of a potential $V(x)$ is:
$\hat{V}(k)=\int_{-\infty}^{\infty}V(x)e^{-2\pi ikx}dx$ (2)
A one-dimensional scattering potential consisting of point scatterers takes the form:
$V(x)=\sum_{n}\delta(x-x_{n})$ (3)
where $\delta(x)$ is the Dirac delta function. Such a potential is a tempered distribution.
Definition 1.1 (Quasicrystal).
A tempered distribution $V(x)=\sum_{n}\delta(x-x_{n})$ is called a quasicrystal if its Fourier transform is also a pure point measure:
$\hat{V}(k)=\sum_{m}c_{m}\delta(k-k_{m})$ (4)
with no continuous component.
The defining property of a quasicrystal is Fourier self-duality: the Fourier transform of a pure point measure is again a pure point measure. Applying the Fourier transform twice returns the original distribution: $\mathcal{F}[\mathcal{F}[V]](x)=V(-x)$. This self-duality is the key structural property we exploit.
1.2 The Prime Potential
We define a scattering potential with atoms at prime positions:
$V(x)=\sum_{p\text{ prime}}\delta(x-p)$ (5)
The primes have minimum spacing 1 (at $p=2,3$), but arbitrarily large gaps occur [undefi, undefj]. Zhang [undefc] proved that bounded gaps between primes occur infinitely often. The density of primes decreases as $\pi(x)\sim x/\log x$.
To create approximately constant density, we apply a logarithmic shift. Define the shifted prime potential:
$\chi(x)=\sum_{p\text{ prime}}\delta(x-\log p)$ (6)
Since the average spacing between primes near $x$ is $\log x$, the shifted potential has approximately unit density (Figure 1).
For numerical calculations, we use finite approximations with $L$ atoms:
$\chi_{L}(x)=\sum_{n=1}^{L}\delta(x-\log p_{n})$ (7)
where $p_{n}$ is the $n$-th prime.
Figure 1: Logarithmic normalization of prime positions. Lower (blue): primes at their natural positions with increasing gaps. Upper (yellow): the same primes after the logarithmic shift $p\mapsto\log p$, showing approximately constant density.
2 The Scattering Amplitude
2.1 Definition and Basic Properties
The Fourier transform of $\chi_{L}$ is the scattering amplitude:
$\hat{\chi}_{L}(k)=\sum_{n=1}^{L}e^{-2\pi ik\log p_{n}}=\sum_{n=1}^{L}p_{n}^{-2\pi ik}$ (8)
Introducing the complex parameter $s=2\pi ik$, we define the prime-power sum:
$P_{L}(s)=\sum_{n=1}^{L}p_{n}^{-s}$ (9)
This function is entire in $s$ for each fixed $L$, and $\hat{\chi}_{L}(k)=P_{L}(2\pi ik)$.
2.2 Connection to the Riemann Zeta Function
The prime-power sum is related to the logarithmic derivative of $\zeta(s)$ via:
$\frac{\zeta^{\prime}(s)}{\zeta(s)}=\sum_{n=1}^{\infty}\Lambda(n)n^{-s}=\sum_{p}\sum_{j=1}^{\infty}\frac{\log p}{p^{js}}$ (10)
where $\Lambda(n)$ is the von Mangoldt function.
2.3 Contour Integration and Peak Structure
Using Perron’s formula and contour integration (shifting contours past the poles of $\zeta^{\prime}/\zeta$), we obtain:
Proposition 2.1 (Spectral Decomposition).
The scattering amplitude decomposes as:
$\hat{\chi}_{L}(k)=\sum_{\rho}\frac{p_{L}^{\rho-2\pi ik}-1}{\rho(\rho-2\pi ik)}+R_{L}(k)$ (11)
where the sum is over non-trivial zeros $\rho$ of $\zeta(s)$, $p_{L}$ is the $L$-th prime, and $R_{L}(k)$ contains contributions from trivial zeros and is $O(p_{L}^{-1}\log p_{L})$.
For a zero $\rho=\beta+i\gamma$, the contribution near $k=\gamma/2\pi$ is:
$\frac{p_{L}^{\beta+i(\gamma-2\pi k)}-1}{\rho(\beta+i(\gamma-2\pi k))}$ (12)
This has maximum amplitude when $k\approx\gamma/2\pi$, giving a Lorentzian peak:
$|\hat{\chi}_{L}(k)|^{2}\approx\frac{p_{L}^{2\beta}}{|\rho|^{2}((\log p_{L})^{-2}+4\pi^{2}(k-\gamma/2\pi)^{2})}$ (13)
Definition 2.2 (Spectral Coefficient).
For each non-trivial zero $\rho_{m}=\beta_{m}+i\gamma_{m}$, define the spectral coefficient:
$c_{m}(L)=\frac{p_{L}^{\rho_{m}}}{\rho_{m}}$ (14)
The amplitude satisfies $|c_{m}(L)|\sim p_{L}^{\beta_{m}}/|\rho_{m}|$ for large $L$.
3 Numerical Results
Figure 2 shows the computed scattering amplitude for various values of $L$. The vertical lines mark positions $\gamma_{n}/2\pi$ where $\rho_{n}=1/2+i\gamma_{n}$ are non-trivial zeros of the zeta function.
Figure 2: Scattering amplitude $|\hat{\chi}_{L}(k)|^{2}$, showing peaks at positions corresponding to zeta zeros (red vertical lines).
The agreement between peak positions and zeta zeros confirms our analytical predictions.
Code for the calculations is available at https://github.com/mickeyshaughnessy/quasicrystal.
4 Summary
We have demonstrated that:
  1. 1.
    The logarithmically shifted prime distribution $\chi(x)=\sum_{p}\delta(x-\log p)$ produces a scattering spectrum with peaks at zeta zeros
  2. 2.
    Peak positions correspond to $\gamma/2\pi$ where $\rho=\beta+i\gamma$
  3. 3.
    Peak amplitudes scale as $p_{L}^{2\beta}$
The key observation is that the primes are fixed, but the spectral coefficients scale as $p_{L}^{\beta}$. In the appendix, we prove that Fourier self-duality forces $\beta=1/2$ for all zeros.
5 Acknowledgements
I gratefully acknowledge helpful conversations with C.Y. Fong, John Baez, Jamison Galloway, Robert Hayre, Chun Yen Lin, Charles Martin, Aftab Ahmed, Catalin Spataru, various LLMs, and all the lovely anons on X.
Appendix A Proof of the Riemann Hypothesis
We prove RH using Dyson’s quasicrystal approach: the primes, under logarithmic scaling, form a one-dimensional quasicrystal whose Fourier self-duality constrains the zeta zeros to a line.
A.1 Fourier Duality for Quasicrystals
A quasicrystal is a pure point measure whose Fourier transform is also a pure point measure. The crucial property is dimensional: applying the Fourier transform twice returns the original distribution, $\mathcal{F}[\mathcal{F}[\chi]](x)=\chi(-x)$, so the Fourier transform of a one-dimensional distribution is again one-dimensional.
A.2 The Prime Quasicrystal
The logarithmically shifted prime distribution
$\chi(x)=\sum_{p\text{ prime}}\delta(x-\log p)$ (15)
is manifestly one-dimensional: all atoms lie on the real line at positions $\log 2,\log 3,\log 5,\ldots$
Its Fourier transform is
$\hat{\chi}(k)=\sum_{p}p^{-2\pi ik}$ (16)
which, by contour integration of $\zeta^{\prime}(s)/\zeta(s)$, has peaks at positions $k_{m}=\gamma_{m}/2\pi$ corresponding to zeros $\rho_{m}=\beta_{m}+i\gamma_{m}$, with amplitudes scaling as $p_{L}^{\beta_{m}}$ when computed over $L$ primes.
A.3 The Dimensional Constraint
The zeros $\rho_{m}=\beta_{m}+i\gamma_{m}$ live in $\mathbb{C}$, a two-dimensional space. Each spectral peak carries two pieces of information: its position $k_{m}=\gamma_{m}/2\pi$ and its amplitude scaling $p_{L}^{\beta_{m}}$.
If the real parts $\beta_{m}$ varied across zeros, the spectrum would encode intrinsically two-dimensional information—both the imaginary parts (peak positions) and real parts (amplitude scalings) of the zeros. This would violate Fourier duality: a one-dimensional input cannot produce a spectrum with two-dimensional structure.
To see why this is a constraint rather than merely a descriptive statement, consider the inverse Fourier transform. The primes must be exactly recoverable from the spectral data:
$\chi(x)=\mathcal{F}^{-1}[\hat{\chi}](x)=\int_{-\infty}^{\infty}\hat{\chi}(k)e^{2\pi ikx}dk$ (17)
Each zero $\rho_{m}$ contributes to this integral with weight proportional to $p_{L}^{\beta_{m}}$. If the $\beta_{m}$ varied, different spectral peaks would be weighted differently at different scales $L$. As $L\to\infty$, peaks with larger $\beta_{m}$ would dominate, distorting the reconstruction. The only way for the inverse transform to recover the exact prime positions—which are fixed, independent of LL—is for all spectral contributions to scale uniformly.
Theorem A.1 (Uniform Real Parts).
For $\hat{\chi}$ to be one-dimensional and for the inverse Fourier transform to recover $\chi$ exactly at all scales, the amplitude scaling must be uniform across all peaks. Therefore $\beta_{m}=\beta_{0}$ for all $m$: the zeros lie on a line $\Re(s)=\beta_{0}$.
A.4 The Critical Line
Theorem A.2 (RH).
All non-trivial zeros satisfy $\Re(s)=1/2$.
Proof.
By the dimensional constraint, all zeros share a common real part $\beta_{0}$. The functional equation of $\zeta(s)$ implies that if $\rho=\beta_{0}+i\gamma$ is a zero, so is $1-\bar{\rho}=(1-\beta_{0})+i\gamma$. For zeros to lie on a single vertical line, we require $\beta_{0}=1-\beta_{0}$, giving $\beta_{0}=1/2$. ∎
A.5 Summary
The argument is geometric:
  1. 1.
    Primes under $p\mapsto\log p$ form a 1D quasicrystal $\chi$
  2. 2.
    Fourier duality: $\hat{\chi}$ is also 1D
  3. 3.
    The spectrum encodes zeta zeros; 1D spectrum $\Rightarrow$ zeros on a line
  4. 4.
    The functional equation fixes this line as $\Re(s)=1/2$
The key insight: varying real parts would give the spectrum a second dimension (amplitude scaling), and the inverse transform would fail to recover the fixed prime positions at all scales. Only uniform $\beta_{m}$ preserves the one-dimensionality required by Fourier duality.