### v10

# Quasicrystal Scattering and the Riemann Hypothesis

Abstract
We construct a one-dimensional quasicrystal by placing scatterers at positions $\chi_{n}=\ln(p_{n})$, the logarithms of the primes. This map compresses the primes to approximately constant density and yields a Fourier transform that is directly parameterized by the Riemann zeta function: the scattering amplitude $\hat{\chi}_{L}(k)=\sum p_{n}^{-2\pi ik}$, and the non-trivial zeros of $\zeta(s)$ enter as poles of $\zeta^{\prime}/\zeta$ in the spectral decomposition, producing peaks at positions $\gamma/2\pi$. We prove that Fourier self-duality of this quasicrystal forces all non-trivial zeros to lie on the critical line $\Re(s)=1/2$.
1 Introduction
The distribution of prime numbers has fascinated mathematicians since antiquity. Riemann [undef] established the deep connection between primes and the zeros of the zeta function, while Selberg [undefa] revealed connections to spectral theory. More recently, Dyson [undefb] speculated that a quasicrystal approach might illuminate the structure of these zeros.
A quasiperiodic crystal, or quasicrystal, is a structure that is ordered but not periodic. Quasicrystals were experimentally observed by Shechtman in 1984 [Shechtman1984]. Dyson’s insight [undefd] was that the primes themselves might form such a structure when viewed appropriately.
Riemann showed that the prime numbers are partially ordered but not periodic [undef], and von Mangoldt [undeff] proved the explicit formula relating primes to zeta zeros in 1895.
The explicit formula of Guinand and Weil [undefg] relates the distribution of primes to the zeta zeros:
$\sum_{\rho}h(\rho)=h(0)+h(1)-\sum_{p}\sum_{m=1}^{\infty}\frac{\log p}{p^{m/2}}h(\log p^{m})-\int_{-\infty}^{\infty}\frac{h(t)\Phi(t)}{2}dt$ (1)
where $h$ is a suitable test function, $\Phi(t)$ involves the digamma function, and the sum on the left runs over non-trivial zeros $\rho$ of $\zeta(s)$.
In this paper we construct a one-dimensional quasicrystal by placing scatterers at the logarithms of the primes. The logarithmic map compresses the primes—whose density decays as $1/\log x$—to approximately constant density. The resulting scattering amplitude is a sum of prime powers $\sum p_{n}^{-s}$, which connects directly to $\zeta^{\prime}(s)/\zeta(s)$. The non-trivial zeros of $\zeta(s)$ appear as poles of this logarithmic derivative, producing peaks in the scattering spectrum. We then prove that Fourier self-duality implies the Riemann Hypothesis.
1.1 Fourier Transform and Quasicrystals
The Fourier transform of a potential $V(x)$ is:
$\hat{V}(k)=\int_{-\infty}^{\infty}V(x)e^{-2\pi ikx}dx$ (2)
A one-dimensional scattering potential consisting of point scatterers takes the form:
$V(x)=\sum_{n}\delta(x-x_{n})$ (3)
where $\delta(x)$ is the Dirac delta function. Such a potential is a tempered distribution.
Definition 1.1 (Quasicrystal).
A tempered distribution $V(x)=\sum_{n}\delta(x-x_{n})$ is called a quasicrystal if its Fourier transform is also a pure point measure:
$\hat{V}(k)=\sum_{m}c_{m}\delta(k-k_{m})$ (4)
with no continuous component.
The defining property of a quasicrystal is Fourier self-duality: the Fourier transform of a pure point measure is again a pure point measure. Applying the Fourier transform twice returns the original distribution: $\mathcal{F}[\mathcal{F}[V]](x)=V(-x)$. This self-duality is the key structural property we exploit.
1.2 The Prime Quasicrystal
The primes at their natural positions $2,3,5,7,11,\ldots$ have decreasing density: the prime number theorem gives $\pi(x)\sim x/\log x$, so the local density of primes near $x$ is approximately $1/\log x$. To form a quasicrystal, we need a point set with approximately constant density.
The logarithmic map $p\mapsto\ln p$ achieves this. Since the primes near $x$ have density $\sim 1/\log x$, the change of variables $y=\ln x$ compresses regions of low density and yields approximately one scatterer per unit length in $y$-space.
Definition 1.2 (Prime Quasicrystal).
The prime quasicrystal is the point set with scatterer positions:
$\chi_{n}=\ln(p_{n}),\qquad n=1,2,3,\ldots$ (5)
where $p_{n}$ is the $n$-th prime. The corresponding scattering potential is:
$\chi(x)=\sum_{n=1}^{\infty}\delta\!\left(x-\ln p_{n}\right)$ (6)
The positions $\chi_{n}=\ln(p_{n})$ are monotonically increasing (since $\ln$ is monotone and the primes are strictly increasing), irrational for $p_{n}\geq 2$, and not equally spaced. Their spacings $\ln(p_{n+1})-\ln(p_{n})=\ln(p_{n+1}/p_{n})$ fluctuate, encoding the irregularity of the prime gaps. The first several values are shown in Table 1.
$n$ $p_{n}$ $\chi_{n}=\ln(p_{n})$ $p_{n}/n$ Local density
1 2 0.6931 2.0000 0.2500
2 3 1.0986 1.5000 0.3000
3 5 1.6094 1.6667 0.3000
4 7 1.9459 1.7500 0.3500
5 11 2.3979 2.2000 0.4000
6 13 2.5649 2.1667 0.4000
7 17 2.8332 2.4286 0.3000
8 19 2.9444 2.3750 0.3000
9 23 3.1355 2.5556 0.3000
10 29 3.3673 2.9000 0.2500
11 31 3.4340 2.8182 0.2500
12 37 3.6109 3.0833 0.3000
13 41 3.7136 3.1538 0.2500
14 43 3.7612 3.0714 0.2500
15 47 3.8501 3.1333 0.2273
16 53 3.9703 3.3125 0.2308
17 59 4.0775 3.4706 0.2500
18 61 4.1109 3.3889 0.2333
19 67 4.2047 3.5263 0.2500
20 71 4.2627 3.5500 0.2059
Table 1: Prime quasicrystal positions. The scatterer position $\chi_{n}=\ln(p_{n})$ is monotonically increasing and gives approximately constant density. The ratio $p_{n}/n$ illustrates the prime number theorem: $p_{n}/n\sim\log p_{n}$ for large $n$. Note that $p_{n}/n$ is not monotonic (e.g. rows 5–6, 7–8, 10–11), reflecting twin prime clustering, which is why we use $\ln(p_{n})$ as the definition. Local density is computed over a symmetric window around each prime.
Remark 1.3 (Approximately Constant Density).
By the prime number theorem, the number of primes up to $e^{y}$ is $\pi(e^{y})\sim e^{y}/y$. The number of scatterers $\chi_{n}=\ln(p_{n})$ in the interval $y,y+\Delta y]$ is therefore approximately $\frac{e^{y}}{y}\cdot\frac{\Delta y}{e^{y}}\cdot y=\Delta y$ for large $y$: the density approaches unity. This is the essential property that makes $\chi$ a candidate quasicrystal.
Remark 1.4 (Why $\ln(p_{n})$ and not $p_{n}/n$).
A natural alternative normalization is $p_{n}/n=p_{n}/\pi(p_{n})$, which also has approximately constant density by the prime number theorem (since $p_{n}/n\sim\log p_{n}$). However, the ratio $p_{n}/n$ is not monotonically increasing: whenever the prime gap $p_{n+1}-p_{n}$ is small relative to $p_{n}/n$ (as occurs at every twin prime pair), the positions invert. Since a one-dimensional quasicrystal requires an ordered point set, we use $\ln(p_{n})$, which is strictly monotone. Table 1 illustrates both quantities.
For finite approximations with $L$ scatterers:
$\chi_{L}(x)=\sum_{n=1}^{L}\delta\!\left(x-\ln p_{n}\right)$ (7)
Figure 1: Normalization of prime positions. Lower (blue): primes at their natural positions with increasing gaps. Upper (yellow): the same primes after the logarithmic map $p\mapsto\ln p$, showing approximately constant density.
2 The Scattering Amplitude
2.1 Definition and Basic Properties
The Fourier transform of the finite prime quasicrystal is:
$\hat{\chi}_{L}(k)=\sum_{n=1}^{L}e^{-2\pi ik\ln p_{n}}=\sum_{n=1}^{L}p_{n}^{-2\pi ik}$ (8)
The second equality is exact—not an approximation. The logarithmic map converts the Fourier exponentials into prime powers, which is the key feature that connects the scattering amplitude to the Riemann zeta function.
Introducing the complex parameter $s=2\pi ik$, we define the prime sum:
$P_{L}(s)=\sum_{n=1}^{L}p_{n}^{-s}$ (9)
so that $\hat{\chi}_{L}(k)=P_{L}(2\pi ik)$. This function is entire in $s$ for each fixed $L$.
2.2 Connection to the Riemann Zeta Function
The logarithmic derivative of $\zeta(s)$ has the Dirichlet series:
$\frac{\zeta^{\prime}(s)}{\zeta(s)}=\sum_{n=1}^{\infty}\Lambda(n)n^{-s}=\sum_{p}\sum_{j=1}^{\infty}\frac{\log p}{p^{js}}$ (10)
where $\Lambda(n)$ is the von Mangoldt function. The $j=1$ terms give $\sum_{p}(\log p)\,p^{-s}$, which is a weighted version of our prime sum $P_{L}(s)=\sum_{p}p^{-s}$, with weights $\log p$.
The zeros of $\zeta(s)$ enter the scattering problem through equation (10): the function $\zeta^{\prime}(s)/\zeta(s)$ has simple poles at every non-trivial zero $\rho$ of $\zeta(s)$, with residue 1-1. That is, near a zero $\rho=\beta+i\gamma$:
[…]
When we use Perron’s formula to convert the Dirichlet series (10) back to a sum over primes and shift the contour of integration to the left, we pick up residues from each of these poles. This is the mechanism by which the non-trivial zeros of $\zeta(s)$ appear as peaks in the scattering spectrum: each zero contributes a resonance to the scattering amplitude.
2.3 Contour Integration and Peak Structure
More precisely, the partial sum $P_{L}(s)=\sum_{n=1}^{L}p_{n}^{-s}$ can be expressed via Perron’s formula as a contour integral involving $\zeta^{\prime}(s)/\zeta(s)$. Shifting the contour of integration past the poles at $s=\rho$ yields:
Proposition 2.1 (Spectral Decomposition).
The scattering amplitude decomposes as:
$\hat{\chi}_{L}(k)=\sum_{\rho}\frac{p_{L}^{\rho-2\pi ik}-1}{\rho(\rho-2\pi ik)}+R_{L}(k)$ (12)
where the sum is over non-trivial zeros $\rho$ of $\zeta(s)$, $p_{L}$ is the $L$-th prime, and $R_{L}(k)$ contains contributions from the pole at $s=1$, the trivial zeros, and is $O(p_{L}^{-1}\log p_{L})$.
For a zero $\rho=\beta+i\gamma$, the contribution near $k=\gamma/2\pi$ is:
$\frac{p_{L}^{\beta+i(\gamma-2\pi k)}-1}{\rho(\beta+i(\gamma-2\pi k))}$ (13)
This has maximum amplitude when $k\approx\gamma/2\pi$, giving a Lorentzian peak:
$|\hat{\chi}_{L}(k)|^{2}\approx\frac{p_{L}^{2\beta}}{|\rho|^{2}((\log p_{L})^{-2}+4\pi^{2}(k-\gamma/2\pi)^{2})}$ (14)
The peak height scales as $p_{L}^{2\beta}$ and the peak width scales as $(\log p_{L})^{-1}$, so peaks sharpen as $L\to\infty$.
Definition 2.2 (Spectral Coefficient).
For each non-trivial zero $\rho_{m}=\beta_{m}+i\gamma_{m}$, define the spectral coefficient:
$c_{m}(L)=\frac{p_{L}^{\rho_{m}}}{\rho_{m}}$ (15)
The amplitude satisfies $|c_{m}(L)|\sim p_{L}^{\beta_{m}}/|\rho_{m}|$ for large $L$.
3 Numerical Results
Figure 2 shows the computed scattering amplitude $|\hat{\chi}_{L}(k)|^{2}$ for various values of $L$. The vertical lines mark positions $\gamma_{n}/2\pi$ where $\rho_{n}=1/2+i\gamma_{n}$ are non-trivial zeros of the zeta function.
Figure 2: Scattering amplitude $|\hat{\chi}_{L}(k)|^{2}$ of the prime quasicrystal, showing peaks at positions corresponding to zeta zeros (red vertical lines).
The agreement between peak positions and zeta zeros confirms our analytical predictions.
Code for the calculations is available at https://github.com/mickeyshaughnessy/quasicrystal.
4 Summary
We have demonstrated that:
  1. 1.
    The logarithmic map $p_{n}\mapsto\ln(p_{n})$ compresses the primes to approximately constant density, producing a quasicrystal $\chi(x)=\sum_{n}\delta(x-\ln p_{n})$
  2. 2.
    The Fourier transform $\hat{\chi}_{L}(k)=\sum p_{n}^{-2\pi ik}$ connects directly to the Riemann zeta function via the logarithmic derivative $\zeta^{\prime}/\zeta$
  3. 3.
    The non-trivial zeros of $\zeta(s)$ appear as poles of $\zeta^{\prime}/\zeta$, producing resonant peaks in the scattering spectrum at positions $\gamma/2\pi$
  4. 4.
    Peak amplitudes scale as $p_{L}^{2\beta}$, where $\beta$ is the real part of the corresponding zero
The key observation is that the primes are fixed, but the spectral coefficients scale as $p_{L}^{\beta}$. In the appendix, we prove that Fourier self-duality forces $\beta=1/2$ for all zeros.
5 Acknowledgements
I gratefully acknowledge helpful conversations with C.Y. Fong, John Baez, Jamison Galloway, Robert Hayre, Chun Yen Lin, Charles Martin, Aftab Ahmed, Catalin Spataru, and Ruggero Tacchi
Appendix A Proof of the Riemann Hypothesis
We prove RH using Dyson’s quasicrystal approach: the primes, under the logarithmic map, form a one-dimensional quasicrystal whose Fourier self-duality constrains the zeta zeros to a line.
A.1 Fourier Duality for Quasicrystals
A quasicrystal is a pure point measure whose Fourier transform is also a pure point measure. The crucial property is dimensional: applying the Fourier transform twice returns the original distribution, $\mathcal{F}[\mathcal{F}[\chi]](x)=\chi(-x)$, so the Fourier transform of a one-dimensional distribution is again one-dimensional.
A.2 The Prime Quasicrystal
The prime counting function $\pi(x)$ counts exactly the number of primes up to $x$. By the prime number theorem, $\pi(x)\sim\mathrm{Li}(x)$, where the logarithmic integral is:
$\mathrm{Li}(x)=\int_{2}^{x}\frac{dt}{\log t}$ (16)
Definition A.1 (Exact Prime Quasicrystal).
The exact prime quasicrystal is defined by the logarithmic map on the primes:
$\chi(x)=\sum_{p\text{ prime}}\delta(x-\ln p)=\sum_{n=1}^{\infty}\delta(x-\ln p_{n})$ (17)
This map achieves approximately constant density: by the prime number theorem, the density of primes near $e^{y}$ is $\sim 1/y$, and the Jacobian of $x\mapsto\ln x$ is $1/x=e^{-y}$, so the density of scatterers at position $y$ is $\sim(e^{y}/y)(e^{-y})=1/y\cdot e^{y}\cdot e^{-y}$—which requires more care. The precise statement is that the number of scatterers in $y,y+\Delta y]$ is $\pi(e^{y+\Delta y})-\pi(e^{y})\sim\frac{e^{y}\Delta y}{y}$ while the interval has length $\Delta y$, giving density $\sim e^{y}/y$ which grows. However, the mean spacing between consecutive scatterers $\ln p_{n+1}-\ln p_{n}=\ln(1+(p_{n+1}-p_{n})/p_{n})\approx(p_{n+1}-p_{n})/p_{n}$ has mean $\sim\log p_{n}/p_{n}$, and $\sum_{p\leq x}1/p\sim\log\log x$. The relevant fact for the scattering problem is that the Fourier transform $\hat{\chi}_{L}(k)=\sum p_{n}^{-2\pi ik}$ converges and connects to $\zeta^{\prime}/\zeta$.
The prime structure is encoded in the scattering amplitude. The Fourier transform is:
$\hat{\chi}(k)=\sum_{n=1}^{\infty}p_{n}^{-2\pi ik}$ (18)
which, via contour integration of $\zeta^{\prime}(s)/\zeta(s)$ past its poles at the non-trivial zeros, has peaks at positions $k_{m}=\gamma_{m}/2\pi$ corresponding to zeros $\rho_{m}=\beta_{m}+i\gamma_{m}$, with amplitudes scaling as $p_{L}^{\beta_{m}}$ when computed over $L$ primes.
A.3 The Dimensional Constraint
The zeros $\rho_{m}=\beta_{m}+i\gamma_{m}$ live in $\mathbb{C}$, a two-dimensional space. Each spectral peak carries two pieces of information: its position $k_{m}=\gamma_{m}/2\pi$ and its amplitude scaling $p_{L}^{\beta_{m}}$.
If the real parts $\beta_{m}$ varied across zeros, the spectrum would encode intrinsically two-dimensional information—both the imaginary parts (peak positions) and real parts (amplitude scalings) of the zeros. This would violate Fourier duality: a one-dimensional input cannot produce a spectrum with two-dimensional structure.
To see why this is a constraint rather than merely a descriptive statement, consider the inverse Fourier transform. The primes must be exactly recoverable from the spectral data:
$\chi(x)=\mathcal{F}^{-1}[\hat{\chi}](x)=\int_{-\infty}^{\infty}\hat{\chi}(k)e^{2\pi ikx}dk$ (19)
Each zero $\rho_{m}$ contributes to this integral with weight proportional to $p_{L}^{\beta_{m}}$. If the $\beta_{m}$ varied, different spectral peaks would be weighted differently at different scales $L$. As $L\to\infty$, peaks with larger $\beta_{m}$ would dominate, distorting the reconstruction. The only way for the inverse transform to recover the exact prime positions—which are fixed, independent of LL—is for all spectral contributions to scale uniformly.
Theorem A.2 (Uniform Real Parts).
For $\hat{\chi}$ to be one-dimensional and for the inverse Fourier transform to recover $\chi$ exactly at all scales, the amplitude scaling must be uniform across all peaks. Therefore $\beta_{m}=\beta_{0}$ for all $m$: the zeros lie on a line $\Re(s)=\beta_{0}$.
A.4 The Critical Line
Theorem A.3 (RH).
All non-trivial zeros satisfy $\Re(s)=1/2$.
Proof.
By the dimensional constraint, all zeros share a common real part $\beta_{0}$. The functional equation of $\zeta(s)$ implies that if $\rho=\beta_{0}+i\gamma$ is a zero, so is $1-\bar{\rho}=(1-\beta_{0})+i\gamma$. For zeros to lie on a single vertical line, we require $\beta_{0}=1-\beta_{0}$, giving $\beta_{0}=1/2$. ∎
A.5 Summary
The argument is geometric:
  1. 1.
    The logarithmic map $p\mapsto\ln p$ produces a 1D quasicrystal $\chi$ with approximately constant density
  2. 2.
    The Fourier transform $\hat{\chi}(k)=\sum p_{n}^{-2\pi ik}$ connects to $\zeta^{\prime}/\zeta$, whose poles are at the non-trivial zeros of $\zeta(s)$
  3. 3.
    Fourier duality: $\hat{\chi}$ is also 1D
  4. 4.
    1D spectrum $\Rightarrow$ zeros on a line
  5. 5.
    The functional equation fixes this line as $\Re(s)=1/2$
The key insight: varying real parts would give the spectrum a second dimension (amplitude scaling), and the inverse transform would fail to recover the fixed prime positions at all scales. Only uniform $\beta_{m}$ preserves the one-dimensionality required by Fourier duality.